OpenStudy (erich):

y=x-1 x-y^2=9 Solve for x, I got the answer of x=the square root of x-8 but it doesnt seem correct

OpenStudy (anonymous):

x-(x-1)^2=9

OpenStudy (anonymous):

x-(x^2-2x+1 = 9

OpenStudy (anonymous):

x-x^2+2x-1=9

OpenStudy (erich):

How did you get 2x in the second step

OpenStudy (anonymous):

(x-1)(x-1) = x^2-2x+1

OpenStudy (erich):

I see

OpenStudy (erich):

so then it becomes -x^2-x+1=9 which means x= the square root of -x-8?

OpenStudy (anonymous):

y=x-1 x-y^2=9 x-(x-1)^2=9 x-x^2+2x-1=9 -x^2+3x-1=9 -x^2+3x-10=0 -1(x^2-3x+10)=0 quadratic equation: x=(3+sqrt(3^2-4(10))/2 x=(3+sqrt(-31))/2 no real solution